OpenStudy (anonymous):

State the domain of the rational function. f(x) = 16/(13-x)

OpenStudy (phi):

the domain are all "valid x values" "non-valid" or "bad" x values are those that cause you to divide by 0, or take the square root of a negative number. In this case, there is only 1 bad x value. can you name it ? you would say the domain is x is all real numbers, except x≠ (fill in the bad x)

OpenStudy (anonymous):

16?

OpenStudy (phi):

what is the denominator of 16/(13-x) when x is 16 ?

OpenStudy (phi):

in other words, what is 13-x when x is 16 ?

OpenStudy (anonymous):

-3

OpenStudy (phi):

and that is not zero, so x=16 is ok. you would get f(16)= 16/-3 which is a number a bad x would cause you to do f(?)= 16/0 and dividing by 0 is bad. The result is undefined. what x would make 13-x zero? you could use algebra and solve for 13-x =0 add x to both sides to solve 13-x+x = 0+x simplify what do you get ?

OpenStudy (anonymous):

13

OpenStudy (phi):

yes, so you want to say the domain for your function is all real numbers, but x ≠ 13

OpenStudy (anonymous):

oh, Okay thanks:)